Mia Breen Corp. produces and sells wind-energy-driven engines. To finance its operations, Mia Breen issued $30,500,000 of 20-year, 10% callable bonds on May 1, Year 1, at their face amount, with interest payable on May 1 and November 1. The fiscal year of the company is the calendar year.

Journalize the entries to record the following selected transactions. Refer to the Chart of Accounts for exact wording of account titles.

Year 1
May 1 Issued the bonds for cash at their face amount.
Nov. 1 Paid the interest on the bonds.
Year 5
Nov. 1 Called the bond issue at 99, the rate provided in the bond indenture. (Omit entry for payment of interest.)
CHART OF ACCOUNTS
Mia Breen Corp.
General Ledger
ASSETS
110 Cash
111 Petty Cash
121 Accounts Receivable
122 Allowance for Doubtful Accounts
126 Interest Receivable
127 Notes Receivable
131 Merchandise Inventory
141 Office Supplies
142 Store Supplies
151 Prepaid Insurance
191 Land
192 Store Equipment
193 Accumulated Depreciation-Store Equipment
194 Office Equipment
195 Accumulated Depreciation-Office Equipment
LIABILITIES
210 Accounts Payable
221 Salaries Payable
231 Sales Tax Payable
232 Interest Payable
241 Notes Payable
251 Bonds Payable
252 Discount on Bonds Payable
253 Premium on Bonds Payable
EQUITY
311 Common Stock
312 Paid-In Capital in Excess of Par-Common Stock
315 Treasury Stock
321 Preferred Stock
322 Paid-In Capital in Excess of Par-Preferred Stock
331 Paid-In Capital from Sale of Treasury Stock
340 Retained Earnings
351 Cash Dividends
352 Stock Dividends
390 Income Summary
REVENUE
410 Sales
610 Interest Revenue
611 Gain on Redemption of Bonds
EXPENSES
510 Cost of Merchandise Sold
515 Credit Card Expense
516 Cash Short and Over
521 Sales Salaries Expense
522 Office Salaries Expense
531 Advertising Expense
532 Delivery Expense
533 Repairs Expense
534 Selling Expenses
535 Rent Expense
536 Insurance Expense
537 Office Supplies Expense
538 Store Supplies Expense
541 Bad Debt Expense
561 Depreciation Expense-Store Equipment
562 Depreciation Expense-Office Equipment
590 Miscellaneous Expense
710 Interest Expense
711 Loss on Redemption of Bonds
Answer #1

Interest calculation for Year 1 Nov = 30,500,000*10%*6/12 = 1,525,000

Gain on redemption = 30,500,000*1% = 305,000

Year 1
May Issued the bonds for cash at face value
Cash Debit          30,500,000
Bonds Payable Credit          30,500,000
Nov Paid the interest on bonds
Interest Expense Debit            1,525,000
Cash Credit            1,525,000
Year 5 Called the bond issue at 99
May Bonds Payable Debit          30,500,000
Gain on redemption of bonds Credit                305,000
Cash Credit          30,195,000